Suppose you have two 8-bit registers, A and B.

A contains the value 11001100

B contains the value 01001111

What is the decimal value of the two's complement binary sum?
Answer #1
A   -   11001100
B   -   01001111
------------------
A+B -(1)00011011
------------------

there was an overflow of 1
00011011 in decimal is 2^4 + 2^3 + 2^1 + 2^0 = 16 + 8 + 2 + 1 = 27

Answer: 27